The length and width of a rectangle are consecutive odd integers. The area of the rectangle is 63 square units. How long is the longest side of the rectangle?
Answer by edjones(8007)

Let the integers be x,and x+2
x(x+2)=63
x^2+2x-63=0
(x+9)(x-7)=0
x=7, x+2=9
.
Ed